# LLMs as Orchestrators: Constraint-Compliant Multi-Agent Optimization for Recommendation Systems

### v1

@@ 1. Introduction
E-commerce recommendation systems play a central role in the web economy and must simultaneously optimize multiple conflicting objectives, including recommendation accuracy, catalog diversity, seller fairness, and new product exposure (Jannach and Adomavicius, 2016; Abdollahpouri et al., 2020). To address these competing goals, prior work has explored weighted objective aggregation and multi-objective optimization techniques (Zheng et al., 2022). However, in most real-world systems, business requirements such as fairness, coverage, and exposure are enforced only as soft penalties, which often leads to persistent constraint violations and undesirable exposure imbalances in deployment (Wang et al., 2024a).
Recent advances in multi-objective recommendation have introduced Pareto-based optimization methods that explicitly model trade-offs among competing objectives (Lin et al., 2024; Ribeiro et al., 2024). While effective at capturing accuracy–diversity trade-offs, most existing approaches treat constraints as additional objectives rather than hard requirements, resulting in Pareto fronts dominated by infeasible solutions when strict business constraints are present (Liang et al., 2024b). This limitation significantly reduces their applicability in production environments where constraint satisfaction is non-negotiable.
@@
As a result, a critical gap remains between multi-objective recommendation research and production requirements. Existing constrained optimization methods either strictly enforce feasibility at the cost of solution quality, or allow infeasible solutions to dominate the Pareto front (Yu et al., 2024; Liang et al., 2024c). At the same time, current LLM-based recommender systems lack mechanisms for adaptive, constraint-aware coordination during optimization. A principled framework that simultaneously achieves strong Pareto performance, hard constraint satisfaction, and adaptive coordination is still missing.
@@ 2.1. Multi-Objective and Constrained Recommendation
Multi-objective recommendation systems aim to optimize multiple competing criteria—such as accuracy, diversity, novelty, and fairness—simultaneously rather than relying on a single objective (Jannach and Adomavicius, 2016). Early approaches typically combine objectives via weighted scalarization, while more recent work adopts Pareto-based optimization to explicitly model trade-offs among objectives (Lin et al., 2024; Zheng et al., 2022). Representative methods include MultiTRON (Ribeiro et al., 2024), which applies Pareto front approximation to session-based recommendation using transformers, and PMORS (Xia et al., 2024), which incorporates temporal forgetting to address preference drift. Industrial systems such as Deep Pareto Reinforcement Learning at Alibaba (Chen et al., 2024) and MOHR at Uber Eats (Luo et al., 2023) further demonstrate the practical value of multi-objective optimization in large-scale marketplaces.
A parallel line of work focuses on fairness, diversity, and exposure control in recommendation systems (Wang et al., 2024a; Li et al., 2024). Prior studies analyze popularity bias and exposure imbalance (Abdollahpouri et al., 2024), propose fairness-aware optimization objectives (Patro et al., 2024; Wu et al., 2024a), and design exploration strategies to improve new-item exposure (Zhang et al., 2024c). While these methods improve long-term ecosystem health, fairness and coverage requirements are typically modeled as soft objectives or regularization terms, providing no guarantees of constraint satisfaction in deployment.
Constrained multi-objective evolutionary optimization (CMOEA) offers principled tools for handling hard constraints (Liang et al., 2024b, c). Techniques such as constraint domination (Deb et al., 2002), $\epsilon$-constraint relaxation (Takahama and Sakai, 2006; Fan et al., 2025), and self-adaptive constraint control (Fan et al., 2019; Liang et al., 2024a) have shown strong performance on benchmark optimization problems. However, their application to recommendation systems remains limited, and existing approaches often either sacrifice solution quality for feasibility or allow infeasible solutions to dominate the Pareto front.
@@ 2.2. LLM Agents for Recommendation and Optimization
Large language models (LLMs) have recently been incorporated into recommendation systems for a wide range of tasks, including item scoring, explanation generation, conversational interaction, and user modeling (Wu et al., 2024b; Liu et al., 2024). Huang et al. (Peng et al., 2025) categorize LLM-based recommender agents into recommender-oriented, interaction-oriented, and simulation-oriented paradigms. Representative systems include Agent4Rec (Zhang et al., 2024a), which uses generative agents with memory for user simulation, MACRec (Wang et al., 2024b), which enables multi-agent collaboration for recommendation, and ToolRec (Zhao et al., 2024), which equips LLMs with external tools for decision-making. Other work applies LLMs to explanation generation (Chen et al., 2023), collaborative filtering (Zhang et al., 2024b), and re-ranking (Hou et al., 2024).
Despite this rapid progress, existing LLM-based recommender systems primarily operate at the level of content understanding or interaction, rather than optimization orchestration. LLMs are typically used to score items, generate natural language outputs, or simulate user behavior, while the underlying optimization processes—such as exploration–exploitation control, constraint handling, and resource allocation—remain rule-based or static. As a result, the potential of LLMs to act as intelligent coordinators for complex, constrained optimization problems in recommendation systems has not been systematically explored.
@@ 4.4. LLM Coordinator Analysis
The LLM coordinator demonstrates adaptive behavior across experiments as shown in Table 5. The average exploitation ratio $\alpha$ trends from 0.6 in early generations to 0.8 in late generations, reflecting a natural shift from exploration to exploitation as the search converges. When feasibility rate drops below 80%, the coordinator increases exploration allocation (reducing $\alpha$ to 0.55), demonstrating constraint-aware adaptation. Qualitatively, the LLM provides interpretable rationales such as: “Increasing exploitation ratio to 0.85 as constraint satisfaction is stable and hypervolume improvement has slowed, suggesting the search is near convergence.” This interpretability aids debugging and provides insights into optimization dynamics.
The coordinator’s decision-making exhibits several notable patterns. In early optimization phases, when constraint violations are common and the Pareto front is poorly approximated, the coordinator favors exploration to discover diverse feasible regions. As optimization progresses and feasibility improves, the coordinator shifts toward exploitation to refine the Pareto front. This adaptive behavior emerges naturally from the LLM’s analysis of optimization metrics without explicit programming of exploration-exploitation schedules.
Coordinator Decision Examples. To concretely illustrate the adaptive coordination, we present three representative decision points from a typical optimization trajectory in the Electronics category. At generation 8, with feasibility rate at 62% and HV at 0.089, the coordinator set $\alpha=0.55$, reasoning: “Feasibility is below target threshold. Prioritizing exploration to discover more constraint-satisfying regions before refining solution quality.” At generation 25, with feasibility at 94% and HV at 0.142, the coordinator increased $\alpha$ to 0.75: “Constraint satisfaction has stabilized. Shifting toward exploitation to improve Pareto front approximation while maintaining feasibility.” At generation 45, detecting HV stagnation (three consecutive generations with <<0.5% improvement), the coordinator temporarily reduced $\alpha$ to 0.60: “Hypervolume progress has plateaued. Increasing exploration to escape potential local optima and discover new trade-off regions.” These examples demonstrate that the LLM coordinator responds not only to absolute metric values but also to optimization dynamics such as convergence patterns and stagnation detection, providing context-aware resource allocation that would require complex hand-crafted rules in traditional approaches.
@@ 5. Discussion
Practical Implications. Our results indicate that constraint strictness is the primary factor affecting recommendation quality in constrained multi-objective optimization. Practitioners should calibrate thresholds based on specific business requirements: stricter constraints for fairness-critical applications, relaxed constraints when diversity and exploration are prioritized. The framework achieves 100% constraint satisfaction across all configurations, making it suitable for production environments where hard constraints are non-negotiable.
@@
Dual-Agent vs. Single-Agent Trade-offs. The dual-agent architecture introduces additional complexity compared to single-population approaches, requiring synchronization between agents and knowledge transfer mechanisms. However, our results demonstrate that this complexity is justified by improved Pareto front coverage and faster convergence. The separation of concerns allows each agent to specialize: the Exploitation Agent maintains high-quality feasible solutions while the Exploration Agent discovers novel trade-offs that would be missed by a unified search strategy.
Deployment Guidelines. Based on our experimental findings, we provide configuration recommendations for different deployment scenarios. For latency-sensitive online recommendations (response time <<500ms), we recommend population size 50, 20 generations, and relaxed constraints ($\theta=0.5$), achieving approximately 5-second optimization time with competitive diversity. For offline batch processing where quality is paramount, population size 200, 100 generations, and strict constraints ($\theta=0.7$) maximize accuracy at the cost of longer runtime (approximately 70 seconds per recommendation set). For balanced production deployments, the default configuration (population 100, 50 generations, $\theta=0.6$) offers a practical compromise. The LLM coordinator should be enabled when interpretability is required for business stakeholders or regulatory compliance, though it can be replaced with rule-based allocation for maximum throughput.
@@
Limitations and Future Work. Our evaluation focuses on three Amazon product categories; broader evaluation across additional domains (e.g., news, music, video) would strengthen generalizability claims. The LLM coordinator adds latency (approximately 2-3 seconds per coordination call) that may be prohibitive for real-time applications requiring sub-100ms response times; distillation to smaller models could address this limitation. Scaling to larger item catalogs (millions of items) requires efficient candidate generation as a preprocessing step, potentially using approximate nearest neighbor search or learned retrieval models. Future work includes real-time deployment with streaming user data, evaluation on larger-scale datasets across multiple domains, integration with production recommendation pipelines, and investigation of different LLM architectures for coordination.
@@ 6. Conclusion
The dual-agent architecture enables specialized optimization through separation of exploration and exploitation, while the LLM coordinator provides intelligent, adaptive resource allocation with interpretable decision rationales. Experiments across three diverse Amazon product categories demonstrate consistent improvements over single-population baselines, with 4.3% relative HV improvement while maintaining full constraint compliance. Ablation studies reveal the critical role of constraint strictness in determining optimization quality, providing practical deployment guidelines for practitioners balancing business constraints with recommendation quality.

### v2

@@ 1. Introduction
E-commerce recommendation systems play a central role in the web economy and must simultaneously optimize multiple conflicting objectives, including recommendation accuracy, catalog diversity, seller fairness, and new product exposure (Jannach and Adomavicius, 2016). To address these competing goals, prior work has explored weighted objective aggregation and multi-objective optimization techniques (Zheng et al., 2022). However, in most real-world systems, business requirements such as fairness, coverage, and exposure are enforced only as soft penalties, which often leads to persistent constraint violations and undesirable exposure imbalances in deployment (Wang et al., 2024).
Recent advances in multi-objective recommendation have introduced Pareto-based optimization methods that explicitly model trade-offs among competing objectives (Lin et al., 2024; Ribeiro et al., 2024). While effective at capturing accuracy–diversity trade-offs, most existing approaches treat constraints as additional objectives rather than hard requirements, resulting in Pareto fronts dominated by infeasible solutions when strict business constraints are present (Liang et al., 2024a). This limitation significantly reduces their applicability in production environments where constraint satisfaction is non-negotiable.
@@
As a result, a critical gap remains between multi-objective recommendation research and production requirements. Existing constrained optimization methods either strictly enforce feasibility at the cost of solution quality, or allow infeasible solutions to dominate the Pareto front (Hao et al., 2024; Liang et al., 2024b). At the same time, current LLM-based recommender systems lack mechanisms for adaptive, constraint-aware coordination during optimization. A principled framework that simultaneously achieves strong Pareto performance, hard constraint satisfaction, and adaptive coordination is still missing.
@@ 2.1. Multi-Objective and Constrained Recommendation
Multi-objective recommendation systems aim to optimize multiple competing criteria—such as accuracy, diversity, novelty, and fairness—simultaneously rather than relying on a single objective (Jannach and Adomavicius, 2016). Early approaches typically combine objectives via weighted scalarization, while more recent work adopts Pareto-based optimization to explicitly model trade-offs among objectives (Lin et al., 2024; Zheng et al., 2022). Representative methods include MultiTRON (Ribeiro et al., 2024), which applies Pareto front approximation to session-based recommendation, and Deep Pareto Reinforcement Learning (Li and Tuzhilin, 2024), which demonstrates practical value in large-scale marketplaces.
A parallel line of work focuses on fairness, diversity, and exposure control in recommendation systems (Wang et al., 2024; Li et al., 2024). Prior studies analyze popularity bias and exposure imbalance (Abdollahpouri et al., 2024), propose fairness-aware optimization objectives (Patro et al., 2024; Wu et al., 2024a), and design exploration strategies to improve new-item exposure (Zhang et al., 2024b). While these methods improve long-term ecosystem health, fairness and coverage requirements are typically modeled as soft objectives or regularization terms, providing no guarantees of constraint satisfaction in deployment.
Constrained multi-objective evolutionary optimization (CMOEA) offers principled tools for handling hard constraints (Liang et al., 2024a, b). Techniques such as constraint domination (Deb et al., 2002), $\epsilon$-constraint relaxation (Takahama and Sakai, 2006), and self-adaptive constraint control (Fan et al., 2019) have shown strong performance on benchmark optimization problems. However, their application to recommendation systems remains limited, and existing approaches often either sacrifice solution quality for feasibility or allow infeasible solutions to dominate the Pareto front.
@@ 2.2. LLM Agents for Recommendation and Optimization
Large language models (LLMs) have recently been incorporated into recommendation systems for a wide range of tasks, including item scoring, explanation generation, conversational interaction, and user modeling (Wu et al., 2024b; Liu et al., 2024). Huang et al. (Peng et al., 2025) categorize LLM-based recommender agents into recommender-oriented, interaction-oriented, and simulation-oriented paradigms. Recent advances include CARTS (Chen et al., 2025), which coordinates multiple LLM agents for structured summarization, ARAG (Maragheh et al., 2025), which combines agentic retrieval with personalized recommendation, GRACE (Ma et al., 2025), which introduces chain-of-thought tokenization, and VL-CLIP (Giahi et al., 2025), which enhances multimodal recommendation. Other work applies LLMs to collaborative filtering (Zhang et al., 2024a) and re-ranking (Hou et al., 2024).
Despite this rapid progress, existing LLM-based recommender systems primarily operate at the level of content understanding or interaction, rather than optimization orchestration. LLMs are typically used to score items, generate natural language outputs, or simulate user behavior, while the underlying optimization processes—such as exploration–exploitation control, constraint handling, and resource allocation—remain rule-based or static (Zhang et al., 2025). Multi-agent frameworks such as CAL-RAG (Forouzandehmehr et al., 2025) demonstrate collaborative agentic reasoning but do not address optimization coordination. As a result, the potential of LLMs to act as intelligent coordinators for complex, constrained optimization problems in recommendation systems has not been systematically explored.
@@ 4.4. LLM Coordinator Analysis
Table 5 shows that $\alpha$ trends from 0.6 (early) to 0.8 (late generations), reflecting a natural exploration-to-exploitation shift. When feasibility drops below 80%, the coordinator increases exploration ($\alpha=0.55$), demonstrating constraint-aware adaptation. The LLM provides interpretable rationales such as: “Increasing exploitation ratio as constraint satisfaction is stable and hypervolume improvement has slowed.” This adaptive behavior emerges naturally from optimization metric analysis without hand-crafted schedules.
Coordinator Decision Examples. At generation 8 (feasibility 62%, HV 0.089), the coordinator set $\alpha=0.55$: “Prioritizing exploration to discover constraint-satisfying regions.” At generation 25 (feasibility 94%, HV 0.142), it increased $\alpha$ to 0.75: “Shifting toward exploitation as constraint satisfaction has stabilized.” At generation 45, detecting HV stagnation, it reduced $\alpha$ to 0.60: “Increasing exploration to escape potential local optima.” These examples demonstrate context-aware resource allocation responding to both metric values and optimization dynamics.
@@ 5. Discussion
Practical Implications. Constraint strictness is the primary factor affecting recommendation quality. Practitioners should calibrate thresholds based on business requirements: stricter for fairness-critical applications, relaxed when diversity is prioritized. The framework achieves 100% constraint satisfaction, suitable for production environments with non-negotiable hard constraints.
@@
Dual-Agent vs. Single-Agent Trade-offs. The dual-agent architecture introduces additional complexity compared to single-population approaches, requiring synchronization and knowledge transfer mechanisms. However, this complexity is justified by improved Pareto front coverage and faster convergence. The Exploitation Agent maintains high-quality feasible solutions while the Exploration Agent discovers novel trade-offs missed by unified search.
Deployment Guidelines. Based on our experimental findings, we provide configuration recommendations for different deployment scenarios. For latency-sensitive online recommendations, we recommend population size 50, 20 generations, and relaxed constraints ($\theta=0.5$), achieving approximately 5-second optimization time. For offline batch processing where quality is paramount, population size 200 with strict constraints ($\theta=0.7$) maximizes accuracy. The default configuration (population 100, 50 generations, $\theta=0.6$) offers a practical compromise. The LLM coordinator should be enabled when interpretability is required for business stakeholders or regulatory compliance.
@@
Limitations and Future Work. Our evaluation focuses on three Amazon product categories; broader evaluation across additional domains (e.g., news, music, video) would strengthen generalizability claims. The LLM coordinator adds latency (approximately 2-3 seconds per coordination call) that may be prohibitive for real-time applications requiring sub-100ms response times; distillation to smaller models could address this limitation. Scaling to larger item catalogs (millions of items) requires efficient candidate generation as a preprocessing step, potentially using approximate nearest neighbor search or learned retrieval models. Future work includes real-time deployment with streaming user data, evaluation on larger-scale datasets across multiple domains, and integration with production recommendation pipelines.
@@ 6. Conclusion
The dual-agent architecture enables specialized optimization through separation of exploration and exploitation, while the LLM coordinator provides intelligent, adaptive resource allocation with interpretable decision rationales. Experiments across three diverse Amazon product categories demonstrate consistent improvements over single-population baselines, with 4.3% relative HV improvement while maintaining full constraint compliance. Ablation studies reveal the critical role of constraint strictness in determining optimization quality, providing practical deployment guidelines for practitioners.